okay well here is my problem i need help on these question and how to do them, these are the direction;
given the coordinate of one endpoint of AB [straight line on top of the AB] and its midpoint M, find the coordinate of the other endpoints
#1 A(0,9) M(2,5)
#2 B(-5,1) M (1,-1)
#3 A(-2,3) M(0.5,0.5)
#4 A (4,2) M (-2,10)
thanks so muchh
PS first time using it lol

Answer by stanbon(48558)

okay well here is my problem i need help on these question and how to do them, these are the direction;
given the coordinate of one endpoint of AB [straight line on top of the AB] and its midpoint M, find the coordinate of the other endpoints (x,y)
#1 A(0,9) M(2,5)
(x+0)/2 = 2 : x = 4
(y+9)/2 = 5 : y = 1
------------------------------
#2 B(-5,1) M (1,-1)
(-5+x)/2 = 1 ; x = 7
(1+y)/2 = -1 ; y = -3
-------------------------------
#3 A(-2,3) M(0.5,0.5)
(-2+x)/2 = 0.5 ; x = 3
(3+y)/2 = 0.5 ; y= -2
-------------------------------
#4 A (4,2) M (-2,10)
(x+4)/2 = -2 ; x = -8
(y+2)/2 = 10 ; y = 18
-------------------------------
Cheers,
Stan H.